How do i simplify and solve this equation 6-2x=5x-9x+10
Answer by richwmiller(17219)

the goal is to get all numbers on one side and all letters (x in this case ) on the other.
We do that by adding to both sides or subtracting from both sides or dividing both sides or multiplying both sides.
but first we combine terms on each side.
so we combine terms on the right side
5x-9x=-4x
so now we have
6-2x=-4x+10
we want all numbers on the right and all letters on the left
so what do we do to remove numbers from the left
we subtract 6 because 6-6=0
so -2x=-4x+4
we want x on the left so we add 4x to both sides because -4x+4x=0
2x=4
divide by 2
x=2
now we check
if the number on the right equals the number on the left then x does = 2
6-2x=5x-9x+10
6-2(2)=5*2-9*2+10
6-4=10-18+10
2=20-18
2=2
ok